Calculations show that a reaction should yield 5.67 g of carbon dioxide gas, CO2 what volume should be expected at 26 C and 782 mmHg


1
Expert's answer
2021-04-27T07:47:51-0400

Mass of CO2 yielded = 5.67 g

Moles of CO2 = "\dfrac{5.67}{44}=0.128\space mol"

Pressure, "P=782\space mm\space of \space Hg=104.258\space kPa"

Temperature, "T=26\degree C=299\space K"

According to ideal gas law,

"PV=nRT"

"V=\dfrac{nRT}{P}=\dfrac{0.128\times8.314\times299}{104258}"

"V=0.00305\space m^3=3.05\space L"